OpenStudy (anonymous):

The dimensions of a rectangular prism are 8 m x 9 m x 10 m. If the dimensions of the rectangular prism are increased by 4, then find the volume of the new prism. a.1284 m3 b.2184 m3 c.4182 m3 d.8142 m3

OpenStudy (dumbcow):

volume = L*w*h just add 4 to each dimension and then multiply those numbers

OpenStudy (anonymous):

i dont get it

OpenStudy (dumbcow):

well whats the volume of a prism of 8 X 9X10 ?

OpenStudy (anonymous):

72

OpenStudy (dumbcow):

make sure you do the third multiplication 8*9=72 then multiply by 10 72*10=720 so volume is 720

OpenStudy (dumbcow):

now do the same thing except add 4 to each number 8+4 = 12 9+4=13 10+4=14

OpenStudy (anonymous):

i still dont understand

OpenStudy (dumbcow):

what do you not understand? volume is length *width*height length=8+4=12 width =9+4=13 height =10+4=14 volume = 12*13*14

OpenStudy (dumbcow):

have you ever used volume before? what class are you in?

OpenStudy (dumbcow):

ok did you figure it out? just multiply the 3 numbers together

OpenStudy (anonymous):

i got it